OpenStudy (anonymous):

Factor (2x-1)^3 - 27 as the difference of two cubes then simplify each fraction.

myininaya (myininaya):

\[(2x-1)^3-3^3=((2x-1)-3)((2x-1)^2+3(2x-1)+3^2)\]

OpenStudy (anonymous):

is that all?

myininaya (myininaya):

-1-3? (2x-1)^2=? 3(2x-1)=? 3^2=?

myininaya (myininaya):

the combine any like terms you might have in the second factor